OpenStudy (anonymous):

find the exact value: arcsin (sqrt(3))/2

OpenStudy (anonymous):

Do you know what value of sin is sqrt{3}/2 ??

OpenStudy (anonymous):

whoops, I mean arcsin (sqrt(3)/2)

OpenStudy (anonymous):

Answer above question..

OpenStudy (anonymous):

\[\sin^{-1} \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

pi/3

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

That's what I thought it was, it just seemed like too easy a question for the final. Thanks.

OpenStudy (anonymous):

\[\sin^{-1}(\frac{\sqrt{3}}{2}) \implies \sin^{-1}(\sin \frac{\pi}{3}) \implies \frac{\pi}{3}\]

OpenStudy (anonymous):

I understand the answer as it relates to the unit circle, but the second operation in that chain confuses me.

OpenStudy (anonymous):

It is okay..

OpenStudy (anonymous):

\[\sin^{-1}(\sin(x)) = x\]

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

You know that sin{pi/3) = sqrt{3}/2, so in second step we are just placing this value in place of sqrt{3}/2..